When the MPC is .75, a decrease in net taxes of $100 billion will increase the equilibrium level of real GDP by

a. $75 billion

b. $100 billion

c. $300 billion

d. $400 billion

Please explain and provide a formula for questions like this if applicable.
Answer #1

Answer

tax Multiplier =-MPC/(1-MPC)

=-0.75/(1-0.75)

=-3

change in GDP=change in taxes * tax multiplier

=-100*(-3) ............. -100 denotes the decrease as the minus sign shows a decrease

=$300 billion

the GDP increases by $300 billion

option c